A change purse contains 14 coins consisting of nickels and dimes and having a total value of $1.10. How many nickels are there?
Answer by stanbon!(97)

Let number of nickels be "x".
Then number of dimes is "14-x"
Value of nickels is (5)x cents
Value of dimes is (10)(14-x) = (140-10x) cents
EQUATION:
Value of nickels + Value of dimes is 110 cents.
5x +140-10x = 110
-5x = -30
x = 6 ( number of nickels)
Cheers,
stan H.